find the area of the shaded region where a is -1.45 and b is 1.87.

write only a number as your answer. round 4 decimal places
Answer #1

Solution: Given that a-1.45 b=1187 Area = Plzea) +P (226) - P(26-145)+ Plz21187) - P(74-145) +[1- P(24187)] = 010735+(1-0.969